A plane flies 435 miles with the wind and 375 miles against the wind in the same length of time. If the speed of the wind is 10 ​mph, find the speed of the plane in still air.
Answer by Boreal(15235)

in x time, the plane flies 435 miles, so speed is 435/x; that speed is p+10
in x time, the plane flies 375 miles, so speed is 375/x;that speed is p-10
p-10=375/x
x=375/p-10
x also = 435/p+10
set those two equal to each other and cross-multiply
375(p+10)=435 (p-10)
375p +3750=435p-4350
60p=8100
p=135 mph. ANSWER
With the wind, the plane flies at 145 mph and covers 435 miles in 3 hours.
Against the wind, the plane flies at 125 mph and covers 375 miles in 3 hours.